A random sample of 300 students from a geographic area indicated that 45 students attended private school. Estimate the true proportion of students attending private schools with 95% confidence


1
Expert's answer
2021-08-05T10:58:09-0400

"n=300 \\

\hat{p}= \frac{45}{300}= 0.15 \\

α=1-0.95=0.05 \\

Z_{α/2}= 1.96 \\

CI= \hat{p}±Z \times \sqrt{ \frac{\hat{p}(1- \hat{p}}{n} } \\

CI = 0.15± 1.96 \times \sqrt{ \frac{0.15(1-0.15)}{300}} \\

= 0.15± 0.0404 \\
( 0.1096 , 0.1904) \\
0.1096<p<0.1904"